# Black hole interiors in the thermal view of scalar-tensor gravity

### v1

@@ Abstract
The thermal view of ST gravity is only defined when $\nabla^{a}\phi$ is timelike and future-oriented and is not applicable when $\nabla^{a}\phi$ and $u^{a}$ are spacelike or null. In first-generation ST gravity, black hole no-hair theorems guarantee that, under suitable conditions, the spacetime geometry outside the horizon of a stationary, asymptotically flat black hole is the same as in GR. The no-hair theorems started with Hawking’s celebrated theorem establishing that stationary black holes that are asymptotically flat in vacuum Brans-Dicke gravity with a massless gravitational scalar field $\phi$ are the same as in GR [31]. Furthermore, the end state of gravitational collapse will produce stationary black holes, which are forced to be axisymmetric [32]. Over the years, the theorem was generalized to include spherical black holes in more general theories [33, 34, 35, 36] and realistic axisymmetric black holes in ST gravities in which the scalar field potential $V(\phi)$ has a minimum [37]. Likewise, the violation of no-hair and the search for black hole hair in alternative theories of gravity is the subject of a large literature (see the recent review [38]).
In theories descrived by the action (1), provided that $V(\phi)$ has a minimum at $\phi_{0}$ and that $\phi$ and $\omega(\phi)$ are not singular outside, or on, the horizon a vacuum static black hole in ST gravity is the same as in GR, i.e., a Kerr black hole because the scalar field is forced to have the constant value $\phi_{0}$ outside the horizon and then gravity reduces to GR. Static, asymptotically flat, spherical black holes in vacuum Brans-Dicke gravity are forced to be Schwarzschild, with line element
@@
Conclusions—We have extended the thermal view of ST gravity to cover the interior region of ST black holes near their spacelike singularities, thus filling an obvious gap in the applications of this new formalism. The discussion of the previous section appears simple, perhaps too simple, but it is a virtue of the new thermal view of ST gravity that it simplifies the description of ST physics and offers a unified frame to interpret technical results in a coherent way. The universality of the Kasner behaviour near spacelike singularities allows one to reach quite general conclusions about the departure of gravity from GR in these regions, where the spacetime structure fails. A first implication is that, if quantization adds a scalar degree of freedom to gravity, as in $f(R)$ gravity [9, 10, 11] or in Starobinski inflation [5], then one should not attempt to quantize the Einstein when looking for Kasner transitions in singularity-free bouncing universes (which are of great current interest), but one should address the field equations of ST gravity instead [50], or adopt an approach not tied to the classical field equations. The conclusion may even be more radical: if the thermal view is taken seriously, gravity and its field equations could be emergent and not suitable for quantization, as proposed in Jacobson’s thermodynamics of spacetime [51, 52]. The latter is quite similar in spirit, although completely different in substance, to the thermal view of ST gravity.

### v2

@@ Abstract
The thermal view of ST gravity is only defined when $\nabla^{a}\phi$ is timelike and future-oriented and is not applicable when $\nabla^{a}\phi$ and $u^{a}$ are spacelike or null. In first-generation ST gravity, black hole no-hair theorems guarantee that, under suitable conditions, the spacetime geometry outside the horizon of a stationary, asymptotically flat black hole is the same as in GR. The no-hair theorems started with Hawking’s celebrated theorem establishing that stationary black holes that are asymptotically flat in vacuum Brans-Dicke gravity with a massless gravitational scalar field $\phi$ are the same as in GR [31]. Furthermore, the end state of gravitational collapse will produce stationary black holes, which are forced to be axisymmetric [32]. Over the years, the theorem was generalized to include spherical black holes in more general theories [33, 34, 35, 36] and realistic axisymmetric black holes in ST gravities in which the scalar field potential $V(\phi)$ has a minimum [37]. Likewise, the violation of no-hair theorems and the search for black hole hair in alternative theories of gravity is the subject of a large literature (see the recent review [38]).
In theories descrived by the action (1), provided that $V(\phi)$ has a minimum at $\phi_{0}$ and that $\phi$ and $\omega(\phi)$ are not singular outside, or on, the horizon a vacuum stationary black hole in ST gravity is the same as in GR, i.e., a Kerr black hole because the scalar field is forced to have the constant value $\phi_{0}$ outside the horizon and then gravity reduces to GR. Static, asymptotically flat, spherical black holes in vacuum Brans-Dicke gravity are forced to be Schwarzschild, with line element
@@
Conclusions—We have extended the thermal view of ST gravity to cover the interior region of ST black holes near their spacelike singularities, thus filling an obvious gap in the applications of this new formalism. The discussion of the previous section appears simple, perhaps too simple, but it is a virtue of the new thermal view of ST gravity that it simplifies the description of ST physics and offers a unified frame to interpret technical results in a coherent way. The universality of the Kasner behaviour near spacelike singularities allows one to reach quite general conclusions about the departure of gravity from GR in these regions, where the spacetime structure fails. A first implication is that, if quantization adds a scalar degree of freedom to gravity, as in $f(R)$ gravity [9, 10, 11] or in Starobinski inflation [5], then one should not attempt to quantize the Einstein equations when looking for Kasner transitions in singularity-free bouncing universes (which are of great current interest), but one should address the field equations of ST gravity instead [50], or adopt an approach not tied to the classical field equations. The conclusion may even be more radical: if the thermal view is taken seriously, gravity and its field equations could be emergent and not suitable for quantization as proposed, e.g., in Jacobson’s thermodynamics of spacetime [51, 52]. The latter is quite similar in spirit, although completely different in substance, to the thermal view of ST gravity.
@@
Singularities are still present in scalar-tensor gravity and the thermal formalism, being purely classical, does nothing to cure them. All that it offers is the realization of how these spacetime singularities are different from those in GR and why—a cure for these singularities will have to come from somewhere else. The understanding of “different” singularities proposed here may have some impact on certain approaches to quantization, but there is no reason to believe this impact to be crucial.
The main point of the new thermal view of scalar-tensor gravity is that it offers a new framework to understand apparently unrelated phenomena that have appeared in the literature over three decades. They are indeed related and the new thermal formalism, which is rather concise, unifies apparently disconnected phenomena in a single view, simplifying their description. However, all these aspects of scalar-tensor gravity need to be revisited one by one in the new picture and then the unifying character of the latter is not manifest. We refer the reader to [19, 20, 21, 22, 23, 24] for details and we will extend the study to new phenomena in future work.